# Aryaman Arora  Neil Rathi  Nikil Roashan Selvam  Róbert Csordás Dan JurafskyChristopher Potts

### v2

@@ Acknowledgements
[…]
Appendix A Formal definitions and details for ATR
For reference, we provide formal definitions for PCFGs and the normal form we use in ATR. [^7] We use similar formalisations of PCFGs as previous work in NLP, e.g. Nowak and Cotterell [2023].
Definition A.1.
  • $\mathcal{N}$ is a finite set of non-terminal symbols;
  • $\Sigma$ is an alphabet of terminal symbols;
  • $\mathrm{S}\in\mathcal{N}$ is a start symbol;
Definition A.2.
A PCFG $\mathcal{G}=\langle\mathcal{N},\Sigma,\mathrm{S},\mathcal{R},p\rangle$ is in Greibach normal form (GNF) if each production rule in $\mathcal{R}$ is of the form $\mathrm{X}\to a~{}\mathrm{X_{1}}~{}\ldots~{}\mathrm{X_{n}}$, where $\mathrm{X}_{1},\ldots,\mathrm{X}_{n}\in\mathcal{N}$ and $n$ may be 00. Similarly, a PCFG is in right-Greibach normal form if each rule is of the form $\mathrm{X}\to\mathrm{X_{1}}~{}\ldots~{}\mathrm{X_{n}}~{}a$.
Definition A.3.
Definition A.4.
A derivation under the PCFG $\mathcal{G}$ is a sequence of strings $[\bm{\alpha}_{0},\ldots,\bm{\alpha}_{m}]$ where $\bm{\alpha}_{0}\in\mathcal{N}$ and each step $\bm{\alpha}_{i+1}$ is formed by a derivation step on $\bm{\alpha}_{i}$. The final string $\bm{\alpha}_{m}\in\Sigma^{*}$ is the yield of the derivation.
A.1 Additional details on ATR
Parent terminals in GNF.
Maximum depth.
[…]
Appendix B Model configurations
@@
short_filter_order $3$
@@
d_conv $4$
@@ Appendix C Task hyperparameters
$\lvert\Sigma\rvert$$\{20,8192\}$
$r_{\Sigma}$ $20$
@@ D.1 Attention needs position embeddings
Due to an initial configuration mistake, we accidentally trained all architectures with absolute position embeddings; in the zoology codebase [Arora et al., 2024a], only Attention is meant to be trained in this way. Fortuitously, this resulted in an interesting ablation: do SSMs, which are usually trained without it, also benefit from position embeddings?
@@
Our results in Figure 7 resoundingly show no: SSMs generally perform worse with position embeddings (PE). Attention is highly dependent on PE; performance on AR drops from $100.00\%$ to $5.62\%$ at $d=256$ with NoPE. Attention lacks recurrence, unlike SSMs, so this is not surprising. However, on ATR, at smaller dimensionalities NoPE actually outperforms PE Attention. Further ablations ought to consider alternative PE methods such as RoPE and Alibi.
D.2 1-layer SSMs learn direct retrieval on AR and ATR
@@
We find comparable performance for direct retrieval models between 1-layer and 2-layer settings on AR (Figure 8(a)). In fact, at $d=256$, 1-layer Mamba ($96.25\%$) outperforms 2-layer Mamba ($91.25\%$), as does Hyena ($31.56\%$ vs. $29.69\%$). 1-layer Based and BaseConv are architecturally identical, so we only report one; that architecture and Attention, both relying on induction in the 2-layer case, fail to learn AR with one layer. On ATR (Figure 8(b)), we see a more noticeable difference with layer count on all architectures, but again Attention, Based, and BaseConv become the worst architectures with one layer (e.g. $96.25\%\to 74.69\%$ for Attention at $d=256$).
D.3 SSMs prefer layer 0 to perform AR
@@
3-layer models perform about the same on AR as 2-layer models across architectures (Figure 9(a)), except for a large drop in performance for Mamba when $d=32$; this may just be an optimisation failure.
@@
For our mechanistic metric, instead of intervening on each block, we intervene at the sequence mixer’s output to the query token in each layer; this tells us if that layer is directly responsible for writing the answer to the output position. We apply the same filter as in section 5.1, with a threshold of $0.01$. Figure 9(b) shows that among performant models, Hyena and Mamba prefer layer 0 for performing AR no matter the layer count; however, some Mamba checkpoints learn the mechanism in the final layer as well (but never layer 1 in a 3-layer model). Attention, Based, and BaseConv prefer layer 1, which is expected since this is the second step of the induction mechanism. However, some checkpoints of Attention and Based also have non-zero attribution score at layer 2 in the 3-layer setting.
D.4 Rightmost sibling queries are trivial for all architectures
@@
Since ATR has hierarchical structure, we attempted an initial experiment with multihop queries; specifically, we present queries where the answer is that terminal’s rightmost sibling terminal. Models are only trained on this type of query, not standard parent queries as reported in the main text. We train with the same settings in section 3.3.
@@
In Figure 10 we show that all models (except Based and BasedConv with $1$ layer, where they only have local convolutions) achieve greater than $80\%$ accuracy at the task at all dimensionalities. We see slight improvement from 1-layer to 2-layer models but at this point performance is saturated and 3-layer does not help. Clearly, this task is extremely simple for all models, even more so than parent queries, and thus does not provide useful signal for comparing architectures.
@@
Parent nodes are guaranteed to be special terminals in our GNF which are sampled from a nonterminal-dependent distribution (see appendix A). However, siblings have a large chance of being fixed terminals specified by the production rule. Additionally, the rightmost sibling of a particular terminal may be itself, if it is the rightmost terminal of its production rule. We speculate that these factors combined make sibling queries easier than parent queries, and thus not a suitable testbed for multihop reasoning.

### v3

@@ 7 Conclusion
Reproducibility statement
We provide our anonymised source code for reproducing all experiments at https://anonymous.4open.science/r/tinylang-1061/. We have specified the hyperparameters for training and evaluation as well as compute usage in appendices˜C and 4.1. We largely use open-source implementations of models from zoology [^7] https://github.com/HazyResearch/zoology (Arora et al., 2024b) as well as the DeltaNet implementation from fla [^8] https://github.com/fla-org/flash-linear-attention, with minor modifications for our mechanistic analysis with the open-source library pyvene [^9] https://github.com/stanfordnlp/pyvene Wu et al. (2024).
@@ Acknowledgements
Appendix A Model configurations
@@
short_filter_order 33
@@
d_conv 44
Appendix B Formalisation of ATR
Param. Description
$H$ Is the head terminal at the left or the right of each production?
$d_{\textrm{max}}$ Maximum depth permitted for the PCFG to generate.
Table 2: Parameters used for constructing a PCFG. We define PCFGs in Greibach Normal Form (GNF); see Appendix B for more details.
For reference, we provide formal definitions for PCFGs and the normal form we use in ATR. [^10] We use similar formalisations of PCFGs as previous work in NLP, e.g. Nowak and Cotterell (2023).
Definition B.1.
  • $\mathcal{N}$ is a finite set of non-terminal symbols;
  • $\Sigma$ is an alphabet of terminal symbols;
  • $\mathrm{S}\in\mathcal{N}$ is a start symbol;
Definition B.2.
A PCFG $\mathcal{G}=\langle\mathcal{N},\Sigma,\mathrm{S},\mathcal{R},p\rangle$ is in Greibach normal form (GNF) if each production rule in $\mathcal{R}$ is of the form $\mathrm{X}\to a~\mathrm{X_{1}}~\ldots~\mathrm{X_{n}}$, where $\mathrm{X}_{1},\ldots,\mathrm{X}_{n}\in\mathcal{N}$ and $n$ may be 0. Similarly, a PCFG is in right-Greibach normal form if each rule is of the form $\mathrm{X}\to\mathrm{X_{1}}~\ldots~\mathrm{X_{n}}~a$.
Definition B.3.
Definition B.4.
A derivation under the PCFG $\mathcal{G}$ is a sequence of strings $\bm{\alpha}_{0},\ldots,\bm{\alpha}_{m}]$ where $\bm{\alpha}_{0}\in\mathcal{N}$ and each step $\bm{\alpha}_{i+1}$ is formed by a derivation step on $\bm{\alpha}_{i}$. The final string $\bm{\alpha}_{m}\in\Sigma^{*}$ is the yield of the derivation.
B.1 Additional details on ATR
Parent terminals in GNF.
Maximum depth.
To enforce maximum depth, we first assign a uniformly random depth score $d:\mathcal{N}\to\mathbb{N}\in\{1,\ldots,\texttt{max\_depth}\}$ to each nonterminal in the vocabulary. Then, for each production rule for each nonterminal $\mathrm{X}$, we only allow nonterminals $\mathrm{Y}$ with $d(\mathrm{Y})>d(\mathrm{X})$ on the right-hand side. Note that this means no recursion is possible.
@@ Appendix C Task hyperparameters
$\lvert\Sigma\rvert$ 2020
$r_{\Sigma}$ 2020
@@ D.1 Attention needs position embeddings
Due to an initial configuration mistake, we accidentally trained all architectures with absolute position embeddings; in the zoology codebase (Arora et al., 2024a), only Attention is meant to be trained in this way. Fortuitously, this resulted in an interesting ablation: do SSMs, which are usually trained without it, also benefit from position embeddings?
@@
Our results in Figure˜9 resoundingly show no: SSMs generally perform worse with position embeddings (PE). Attention is highly dependent on PE; performance on AR drops from $100.00\%$ to $5.62\%$ at $d=256$ with NoPE. Attention lacks recurrence, unlike SSMs, so this is not surprising. However, on ATR, at smaller dimensionalities NoPE actually outperforms PE Attention. Further ablations ought to consider alternative PE methods such as RoPE and Alibi.
D.2 Mamba’s solution to ATR does generalise
We reuse the settings from our ATR experiment ($L=5,\lvert\Sigma\rvert=20$) and construct a new dataset with a train–test split on query–answer pairs. Specifically, $80\%$ of possible unique query–answer pairs are provided in the training set, while $20\%$ are only in the test set and thus never trained on. We seek to assess whether models learn a general mechanism for parent–child relations in ATR or if the impressive results of Mamba (as well as Attention and Based) are merely the result of better memorisation of the PCFG parameters. This setup is akin to Wang et al. (2024)’s technique of train–test split on multi-hop queries; we provide supervision on individual query and answer types, but not on some compositions of them.
Behavioural results.
We select the checkpoint with the highest dev accuracy for each architectural and dimensionality setting, after sweeping LR. We plot the dev and test accuracies of each of these checkpoints in Figure˜10(a); all models have much lower test accuracy (e.g. Attention with $d=256$ has $95.62\%$ dev and $68.12\%$ test accuracy). Attention achieves the greatest dev accuracies on $d\geq 32$. Mamba’s relative ranking is lower than on the in-distribution setting in section˜5.2, but it still achieves the overall second-highest dev accuracy ($65.00\%$ at $d=128$). Surprisingly, H3 generalises well despite its poor dev accuracy, beating Mamba on test accuracy in $3$ out of $5$ settings.
We compare dev and test accuracies across all LRs in Figure˜10(b). We find that while Mamba does have unusually high dev accuracy given a selected test accuracy (indicating greater memorisation than models with other architectures), its dev accuracy is still generally higher than non-Attention architectures. Interestingly, H3 has nearly Attention-level generalisation while BaseConv exhibits vanishingly little generalisation. Overall, behavioural metrics show that Mamba does nontrivially generalise on ATR, albeit not as well as Attention.
Mechanistic analysis.
We report a summary of attribution scores at different tokens (key, query, value), comparing on dev and test sets across all checkpoints in Figure˜10(c). We find largely consistent mechanisms underlying behaviour on both dev and test, and these match attribution scores on ATR without train–test split. The only exception is that BasedConv does induction on the dev set but not nearly as much on the test set; its induction mechanism is more brittle than Attention and Based.
D.3 1-layer SSMs learn direct retrieval on AR and ATR
@@ D.2 Mamba’s solution to ATR does generalise
We find comparable performance for direct retrieval models between 1-layer and 2-layer settings on AR (Figure˜11(a)). In fact, at $d=256$, 1-layer Mamba ($96.25\%$) outperforms 2-layer Mamba ($91.25\%$), as does Hyena ($31.56\%$ vs. $29.69\%$). 1-layer Based and BaseConv are architecturally identical, so we only report one; that architecture and Attention, both relying on induction in the 2-layer case, fail to learn AR with one layer. On ATR (Figure˜11(b)), we see a more noticeable difference with layer count on all architectures, but again Attention, Based, and BaseConv become the worst architectures with one layer (e.g. $96.25\%\to 74.69\%$ for Attention at $d=256$).
D.4 SSMs prefer layer 0 to perform AR
@@
3-layer models perform about the same on AR as 2-layer models across architectures (Figure˜12(a)), except for a large drop in performance for Mamba when $d=32$; this may just be an optimisation failure.
@@
For our mechanistic metric, instead of intervening on each block, we intervene at the sequence mixer’s output to the query token in each layer; this tells us if that layer is directly responsible for writing the answer to the output position. We apply the same filter as in section˜4.2, with a threshold of $0.01$. Figure˜12(b) shows that among performant models, Hyena and Mamba prefer layer 0 for performing AR no matter the layer count; however, some Mamba checkpoints learn the mechanism in the final layer as well (but never layer 1 in a 3-layer model). Attention, Based, and BaseConv prefer layer 1, which is expected since this is the second step of the induction mechanism. However, some checkpoints of Attention and Based also have non-zero attribution score at layer 2 in the 3-layer setting.
D.5 Rightmost sibling queries are trivial for all architectures
@@
Since ATR has hierarchical structure, we attempted an initial experiment with multihop queries; specifically, we present queries where the answer is that terminal’s rightmost sibling terminal. Models are only trained on this type of query, not standard parent queries as reported in the main text. We train with the same settings in section˜4.1.
@@
In Figure˜13 we show that all models (except Based and BasedConv with $1$ layer, where they only have local convolutions) achieve greater than $80\%$ accuracy at the task at all dimensionalities. We see slight improvement from 1-layer to 2-layer models but at this point performance is saturated and 3-layer does not help. Clearly, this task is extremely simple for all models, even more so than parent queries, and thus does not provide useful signal for comparing architectures.
@@
Parent nodes are guaranteed to be special terminals in our GNF which are sampled from a nonterminal-dependent distribution (see appendix˜B). However, siblings have a large chance of being fixed terminals specified by the production rule. Additionally, the rightmost sibling of a particular terminal may be itself, if it is the rightmost terminal of its production rule. We speculate that these factors combined make sibling queries easier than parent queries, and thus not a suitable testbed for multihop reasoning.